OpenStudy (anonymous):

Find the exact value by using a half-angle identity. cosine of negative pi divided by eight

OpenStudy (anonymous):

sorry i dont understand what you mean by that could you please reiterate @SolomonZelman

OpenStudy (solomonzelman):

` π = 1 8 0 º `

OpenStudy (solomonzelman):

Then divide 180 by -8, so you get -22.5 And that is equivalent to (1/2) × 45

OpenStudy (anonymous):

yes i understand that if you could how does that contribute to the problem @SolomonZelman

OpenStudy (anonymous):

is that using the half angle identity though? @SolomonZelman

OpenStudy (anonymous):

if possible could anyone come and help me through this @ParthKohli @jim_thompson5910

OpenStudy (anonymous):

From Mathematica 9 Home Edition:\[\text{ToRadicals}\left[\cos \left(-\frac{\pi }{8}\right)\right]=\frac{\sqrt{2+\sqrt{2}}}{2} \]

OpenStudy (anonymous):

thank you for the answer but could you help me go through the problem if you could

OpenStudy (anonymous):

@robtobey